∫ dx / (x 2 (1+x 2 ) 1/2 =................................................?

∫ dx / (x2(1+x2)1/2=................................................?

Grade:12

1 Answers

Khimraj
6 years ago
I =∫ dx / (x2(1+x2)1/2)
I = ∫ dx / (x3(1/x2  +  1 )1/2 )
put (1/x2) + 1 = t
(-2/x3)dx = dt
(1/x3)dx = -dt/2
I = ∫(1/t1/2)(-dt/2)
I = -1/2∫dt/t1/2
I = –t1/2 = – [ (1/x2) + 1]
Hope it clears. If you like answer then please approve it.